OpenStudy (anonymous):

The quotient of a number and 2 is the same as the difference of the number doubled and 3. What is the number?

OpenStudy (cwrw238):

x/2 = 2x - 3 solve for x

OpenStudy (anonymous):

let x = the number x/2 = 2x - 3 x = 2(2x -3) x = 4x - 6 0 = 3x - 6 3x = 6 x = 2